OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2.

OpenStudy (anonymous):

y=mx+b is slope intercept where m is your slope and your y intercept is when x = 0 (so the B value)

OpenStudy (anonymous):

The line of equation in slope intercept form is given as: y=mx+c, where m=slope =3 and c= y-intercept-2 i.e. y= 3x-2 is the required eq.